OpenStudy (anonymous):

Write an equation for each translation of y=[x] 10.5 units left

OpenStudy (mathmate):

The general equation for translating a function f(x) by (h,k) is T(h,k) f(x)= f(x-h)+k where T(h,k) is the translation operator. For example, if f(x)=\(x^2+4x+2\), then T(5,2) f(x)=\((x-5)^2+4(x-5)+2=x^2-6x+7\) will translate f(x) 5 units to the right, and 2 units upwards. The important to note is that x is replaced by (x-h) wherever x appears, and do not forget the parentheses. By the way, [x] represents the step function as shown below, but the translation works as described above. 

OpenStudy (anonymous):

so my answer would be y = | x – 10.5 |